OpenStudy (anonymous):

mrs jennings spent 14 1/6 hours at the gym last month. mr jennings spent 12 5/6 hours at the gym last month. how much more time did mrs jennings spend at the gym than mr jennings

OpenStudy (anonymous):

You have to subtract mr. jennings hours from mrs. jennings hours

OpenStudy (anonymous):

so first do 14 - 12 to get 2, then do 1/6-5/6, which would equal (-4/6), so now oyu have to take 4/6 away from 2, to get your answer

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

not exactly, 2 - 4/6, well lets make 2 into 12/6, since thats the equivalent, now do 12/6 - 4/6

OpenStudy (anonymous):

8/6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

carson has only $20 bills and $10 bills in her wallet. the total value of the bills is $50. She has 1 more $20 bill then $10 bills. How many of each kind of bill does carson have?

OpenStudy (anonymous):

hey joshua u there?

OpenStudy (anonymous):

ello

OpenStudy (anonymous):

she has two twentys and a ten

OpenStudy (anonymous):

forgive me for not replying earlier, i'm cooking my lunches for the week and trying to do my calc all at the same time lol

OpenStudy (anonymous):

@sarap29 still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

carson has only $20 bills and $10 bills in her wallet. the total value of the bills is $50. She has 1 more $20 bill then $10 bills. How many of each kind of bill does carson have?

OpenStudy (anonymous):

10x +20(x+1)=50?

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

she has two twenties and one ten

OpenStudy (anonymous):

10x+20x+20=50 10x+20x=30 30x=30 x=1

OpenStudy (anonymous):

question 2) Sara bought shirts an hats on sale. the shirt cost $12 each and the hats cost $9 each . she bought 4 more than hats. if she spent a total of $111, how many shirts an hats did she buy/

OpenStudy (anonymous):

thank u for question 1 :)

OpenStudy (anonymous):

question 3) kay looked in her closet to find that 7/10 of her shirts were black an 1/10 of her shirts are white. what fraction of her shirts are black an white?

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

hey yeah

OpenStudy (anonymous):

ok just checking :)

OpenStudy (anonymous):

she bought 7 shirts and 3 hats

OpenStudy (anonymous):

i dont understand question 3 but i think its 8/10

OpenStudy (anonymous):

whats question 2?

OpenStudy (anonymous):

question 4) whats the least common multiple of 252

OpenStudy (anonymous):

i dunno lol

OpenStudy (anonymous):

least common multiple of 252 and what

OpenStudy (anonymous):

kevin is taking a poll of hi gym class to find their favorite sport he finds 2/9 of the class prefers soccer, while another 2/3 prefers baseball. what fraction of the class prefers either soccer or baseball?

OpenStudy (anonymous):

question 6) mr garrett is building a sidewalk. he mixed 5/8 tons of sand an 1/4 ton of concrete . how much did the mixture weigh?

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

for the gym class question, the answer is 8/9

OpenStudy (anonymous):

1/4=2/8

OpenStudy (anonymous):

so for the concrete question, it's 5/8+2/8, can you solve that?